# Entropy-driven entanglement forging

### v1

@@ Appendix B Further details on the convergence of EDEF in the nuclear shell model
Finally, in the main text we briefly mentioned how results of 28Ne compare to those of 26Ne, as an example with less particle/hole symmetry than the simulations in the main text, but for which the first layer of entanglement forging also performs well. We include here such simulation for completeness.
Figure 15: $\epsilon_{E}$ (top plot) and $I$ (bottom plot) for 26Ne as a function of the number iterations for 0 and 1 cuts.
@@
We plot the same errors as in section IV.2 for 0- and 1-cuts simulations for 26Ne in Fig. 15 with solid red and dashed blue lines, respectively. In this case both lines have a very similar behaviour, decreasing gradually down to $\epsilon_{E}=0.013$, $I=0.11$, for 0 cuts, and to $\epsilon_{E}=0.016$, $I=0.16$ for 1 cut. Neither ADAPT-VQE has converged into a local minimum at the 100th iteration, and in the case of of 1-cut ADAPT-VQE, the infidelity and errors are still well above the values obtained from the decomposition of the exact solution into six product states, $\epsilon_{E}^{(1)}=0.011$ and $I^{(1)}=0.044$.
Appendix C Optimization improvements
@@
The implementation of a variational algorithm such as ADAPT-VQE can fall short even when the ansatz and operator pool are adequate for the physical problem due to the necessity of training the parameters. Different implementations of the gradient descent work better for specific problems, but in general they will all run into resource problems as we try to deal with bigger systems. In this section, we present some optimization techniques that we implemented to increase the size of simulatable systems given a specific resource budget, and were specially useful to extend the simulation of the 60Ti nucleus into more layers.
@@
Firstly, the more expensive step is often the parameter tuning. However, the parameters found after the update are in general quite close to the previous step. Instead of optimizing over all parameters each time, one can find the optimal parameter only for the last operator that was added, and perform a complete update of all parameters every $l$ rounds. Because now the operator with the biggest gradient is found with different parameters, the list of chosen operators may vary from the usual approach (equivalent to $l\neq 1$), making $l$ an hyperparameter $l$ that needs tuning. Despite this, we found that the precision of the training after the same amount of steps was close with high probability, even with a different set of chosen operators – this is consistent with the fact that the same ground state can be reached with different compositions of unitaries, due to non-uniqueness of unitary decomposition. While performance increases with $l$, one cannot make it too large because with fixed parameters, even if close to the optimal value, the error can only be decreased by adding extra operators, and a deep circuit becomes exponentially expensive to simulate.
@@
In our entanglement forging approach to ADAPT-VQE, we find that the main circuit, the one with the leading coefficient on the Schmidt decomposition, is often prioritized by the operator choice. As we have mentioned, deep circuits are harder to simulate, so one can exhaust the resources available by always choosing the optimal operator even if some of the subleading circuits are almost empty. If we sometimes choose to place an operator in a circuit that is shallow (and therefore easy to compute), we can improve our error with a relatively inexpensive step (specifically, relative to the progress in the algorithm). In addition, since gradients are evaluated at a fixed point and the optimization happens over all parameter space, the error can decrease more on a single step by adding operators that have a non-optimal gradient. This makes this technique have an even bigger impact, as some of these steps will be more efficient and also more effective. We implemented this idea by excluding the “full” circuit from the choice of operators. This was done with an ad-hoc heuristic based on the optimisation duration in the specific setup used, but can also be generalized into an hyperparameter $\varphi$.

### v2

@@ Appendix C Results for a different isotope
In the main text, we briefly mentioned how results of 28Ne compare to those of 26Ne. This isotope is an example with less particle/hole symmetry than the simulations in the main text, but for which the first layer of entanglement forging also performs well. We include here the details of such simulation for completeness.
@@
We summarize the EDEF results for 26Ne in Fig. 15. The top panel shows the relative energy energy as a function of the iteration number, in line with the results presented in Fig. 10 for 28Ne. We plot results for 0- and 1-cut simulations with solid red and dashed blue lines, respectively. In this case, both lines have a very similar behaviour, decreasing gradually down to $\epsilon_{E}=0.013$ for full ADAPT-VQE and to $\epsilon_{E}=0.016$ for the 1-cut case. In terms of energy, ADAPT-VQE has not converged into a local minimum at the $100^{\textrm{th}}$ iteration. The 1-cut simulation follows a very similar trend and has not yet reached the bound associated to the Schmidt decomposition.
The bottom panels focus on the evolution of the infidelity for the two methods. These should be compared to Fig. 7 in the main text. ADAPT-VQE here performs rather well, and the infidelity with respect to the ground state improves quickly between iterations $15$ and $20$, followed by a steady decrease afterwards. The infidelity reaches a value of $I=0.11$ at iteration $100$. In contrast, the evolution of the infidelity for the 1-cut simulation is somewhat more erratic. After an initial decrease, the infidelity increases between iterations $10$ and $30$, only to decrease later on at a rate that is more or less commensurate with the ADAPT-VQE results. The infidelity at the final step is $I=0.16$, a value that is still well above the corresponding Schmidt decomposition, $I^{(1)}=0.044$. Overall, the two pictures indicate that more iterations are required to reach a good representation of the ground state of 26Ne. The 1-cut EDEF approach may be able to reach this state with far fewer resources than the associated ADAPT-VQE simulation.
Figure 15: $\epsilon_{E}$ (top) and $I$ (bottom panel) for 26Ne as a function of the number iterations for ADAPT-VQE simulation (solid line) and 1-cut EDEF.
Appendix D Optimization improvements
@@
The implementation of a variational algorithm such as ADAPT-VQE can fall short even when the ansatz and operator pool are adequate for the physical problem due to the necessity of training the parameters. Different implementations of the gradient descent work better for specific problems, but in general they may all run into resource problems as we try to deal with bigger systems. In this section, we present some optimization techniques that we implemented to increase the size of simulatable systems given a specific resource budget, which were specially useful to extend the simulation of the 60Ti nucleus into more layers.
@@
First, we note that the more expensive step in our simulations is often the parameter tuning. However, we have heuristically found that the parameters after an update $k$ are in general quite close to those of the previous step, $k-1$. Instead of optimizing over all parameters each time, one can find the optimal parameter only for the last operator that was added, $\theta_{k}$. We can then perform a complete update of all the parameters every $l$ rounds. Because now the operator with the biggest gradient is found with different parameters than the standard minimisation (equivalent to $l=1$), the list of chosen operators may change from the usual approach. This turns $l$ into a hyperparameter that may require tuning. In spite of these differences, we find that the precision of the training after the same number of steps is close with high probability, with very little dependence on the value of $l$. This is true even when the minimization employs a different set of chosen operators, consistent with the fact that the same ground state can be reached with different compositions of unitaries (due to the non-uniqueness of the unitary decomposition). While in general performance increases with $l$, one cannot make it too large because with fixed parameters the error can only be decreased by adding extra operators, and a deep circuit becomes exponentially expensive to simulate. Here, we choose $l=10$ for our 60Ti simulations.
@@
In our entanglement forging approach to ADAPT-VQE, we find that the main circuit, the one with the leading coefficient on the Schmidt decomposition, is often prioritized by the operator choice. As we have mentioned, deep circuits are harder to simulate, so one can exhaust the resources available by always choosing the optimal operator even if some of the subleading circuits are almost empty. If we sometimes choose to place an operator in a circuit that is shallow (and therefore easy to compute), we can improve our error with a relatively inexpensive step (specifically, relative to the progress in the algorithm). In addition, since gradients are evaluated at a fixed point and the optimization happens over all parameter space, the error can decrease more on a single step by adding operators that have a non-optimal gradient. This makes this technique have an even bigger impact, as some of these steps will be more efficient and also more effective. We implemented this idea by excluding the “full” circuit from the choice of operators. While this was done with an ad-hoc heuristic based on the optimisation duration in our specific setup, the approach could potentially be generalized by introducing a new hyperparameter $\varphi$.